When a group of students sits in rows of six
chairs, the last row has one empty chair. When
this same group sits in rows with seven chairs,
there is once again one empty chair. What is the
least number of students in this group?

Answer by stanbon(75887)

When a group of students sits in rows of six
chairs, the last row has one empty chair. When
this same group sits in rows with seven chairs,
there is once again one empty chair. What is the
least number of students in this group?
----------
y1 = (x+1)/6
y2 = (x+1)/7
-----
Find positive integer of the equation::
Ans: x = 41
-------------
Note:
(41+1)/6 = 7
(41+1)/7 = 6
--------------------
Cheers,
Stan H.
-----------------